# LLAMAFUZZ: Large Language Model Enhanced Greybox Fuzzing

### v3

@@ 4.2 Implementation details
To evaluate the potential of LLMs in addressing the limitations of traditional fuzzing for structured data, we implemented LLAMAFUZZ by extending AFL++. We used llama2-7b-chat-hf Touvron et al. (2023) as the base model, which was powerful and efficient. Since LLAMAFUZZ was built on top of AFL++, any observed difference can be attributed to our changes to the implementation of LLM mutation. To prevent seed memorization, we excluded all experimental seeds from the fine-tuning dataset.
We followed the standard instructions provided by the benchmark developers to build the fuzzing targets Serebryany (2017); Metzman et al. (2021); Hazimeh et al. (2020). For real-world programs tested under OSS-Fuzz, we utilized the default initial seed corpus as outlined by OSS-FUZZ Serebryany (2017). Similarly, during our experiments with the Magma benchmark, we selected default initial seeds specified by the benchmark developers. More details, source code, and related artifacts are provided in AppendixA.4.
@@
Figure 3: Distribution of the average number of bugs triggered over 24 hours.
@@
Figure 4: Arithmetic mean number of bugs identified for each project per trial. The black line denotes the 95% confidence interval.
4.3 LLAMAFUZZ find more bugs
@@
Table 2: Average branch coverage achieved by our approach (LLAMAFUZZ) and the baseline (AFL++). We report the average branch coverage, p-values produced by the Mann-Whitney U Test, and the Vargha-Delaney statistics ($\hat{A}_{12}$). For assessing effect size, we use the labels -, M, and L to represent negligible, medium, and large effects, respectively.
@@
binutils fuzz_nm 13 96913\,969 $9017$ 54.9154.91% <0.01 L(1)
fuzz_objcopy 22 11822\,118 12 49412\,494 77.03% <0.01 L(1)
fuzz_readelf $6552$ $4437$ 47.67% <0.01 L(1)
fuzz_strings $6441$ $5295$ 21.64% <0.01 L(1)
bloaty fuzz_target $5972$ $5722$ 4.37% <0.01 L(1)
freetype2 freetype2-ftfuzzer 10 52110\,521 $9978$ 5.45% <0.01 L(1)
grok grk_decompress_fuzzer $3721$ $2313$ 60.87% <0.01 L(1)
@@
4.4 Performance comparison on OSS-Fuzz
@@
First, we evaluated the performance of LLAMAFUZZ against specialized grammar-based fuzzers, including Gramatron Srivastava and Payer (2021), Grimoire Blazytko et al. (2019), and Nautilus Aschermann et al. (2019). We selected a common set of fuzzing targets compatible with all specialized grammar-based fuzzers, including mruby, PHP, and quickjs. Next, we conducted a more comprehensive evaluation using real-world programs. Given that grammar-based fuzzers are limited to well-defined, specialized targets, AFL++ was chosen as the reference competitor, representing the state-of-the-art in greybox fuzzing subsection 4.3.
Figure 5: The distribution of reached bug achieved by LLAMAFUZZ, AFL++, Gramatron, Grimoire, and Nautilus.
Table 3: Average branch coverage achieved by LLAMAFUZZ, baseline with default initial seeds (AFL++) and baseline with default initial seeds combined with trimmed seeds set from LLAMAFUZZ’s fine-tuning dataset (AFL++).
Fuzz target Fuzz object Branch coverage (avg)
LLAMAFUZZ AFL++ AFL++ Improv. p-value $\hat{A}_{12}$
binutils fuzz_nm 13 96913\,969 $9017$ 10 98310\,983 27.18% <0.01 L(1)
fuzz_objcopy 22 11822\,118 12 49412\,494 13 57513\,575 62.93% <0.01 L(1)
fuzz_readelf $6552$ $4437$ $5789$ 13.79% <0.01 L(1)
fuzz_strings $6441$ $5295$ $5220$ 23.39% <0.01 L(1)
bloaty fuzz_target $5972$ $5722$ $5918$ 0.90% 0.37 M(0.7)
grok grk_decompress_fuzzer $3721$ $2313$ $2629$ 41.54% 0.02 L(1)
kamailio fuzz_parse_msg $3743$ $2692$ $2521$ 48.49% <0.01 L(1)
libavc avc_dec_fuzzer $9872$ $9838$ $9847$ 0.26% 0.01 L(1)
@@
5 Analysis
@@
5.1 Did the LLM improve fuzzing by memorizing the fine-tuning data?
@@
Figure 6: Code coverage growth of LLAMAFUZZ and AFL++ on the binutils-nm target over time. Black triangles mark LLM-generated seeds, while the red vertical line highlights seeds directly sourced from LLM. The green line represents LLAMAFUZZ’s coverage, and the grey background indicates AFL++’s coverage range.
5.2 How did LLM augment the fuzzing process?
5.3 Case study
Table 4: Byte-level comparison of an original PNG seed and its AFL- and LLAMAFUZZ-mutated variants.
@@ 5.3 Case study
6 Conclusion
7 Limitation
As aforementioned, our dual-layer structure allows easy replacement of different LLMs, including both open-source and closed-source models. We present an ablation study in AppendixB.2 to evaluate the generalizability and effectiveness of our approach across different LLMs, including llama2-7b-chat-hf Touvron et al. (2023), LLaMA-3-8B-InstructDubey et al. (2024), Mistral-7B-Instruct-v0.2Jiang et al. (2023), and Qwen2-7B-InstructYang et al. (2024a). A full re-run of every experiment on every model is left to future work as resources permit, but our representative trials already demonstrate that the proposed method consistently transfers across architectures.
Appendix A Implementation details
A.1 Fine-tuning
A.2 Asynchronous Queue Design
As summarized in Algorithm 1, our hybrid fuzzer orchestrates three concurrent components—Main Fuzzing Loop, LLM Mutation Thread, and AFL Mutation Thread. This design decouples resource‐intensive, semantics‐driven mutations from the high‐throughput fuzzing loop, enabling non‐blocking operation and easy scaling or replacement of the LLM backend.
Main Fuzzing Loop
LLM Mutation Thread
AFL Mutation Thread
Algorithm 1 Hybrid Fuzzer with LLM- and AFL-based Mutations
1:Initialization: start_fuzzer(); llm_mutate_thread(); afl_mutate_thread()
2:function start_fuzzer
3:  while True do
4:   if Queue_LLM is not empty then
5:     seed $\leftarrow$ dequeue(Queue_LLM)
6:   else
7:     seed $\leftarrow$ dequeue(Queue_AFL)
8:   end if
9:   if run_fuzzer(seed) yields an interesting result then
10:     enqueue(Queue_LLM_unmutated, seed)
11:     enqueue(Queue_AFL_unmutated, seed)
12:   end if
13:  end while
14:end function
15:function llm_mutate_thread
16:  while True do
17:   if not Queue_LLM_unmutated.empty() then
18:     seed $\leftarrow$ dequeue(Queue_LLM_unmutated)
19:     hex_seed $\leftarrow$ convert_to_hex(seed)
20:     mutated_seed $\leftarrow$ LLM_generate(hex_seed) $\triangleright$ LLM-based mutation on GPU
21:     enqueue(Queue_LLM, mutated_seed)
22:   end if
23:  end while
24:end function
25:function afl_mutate_thread
26:  while True do
27:   if not Queue_AFL_unmutated.empty() then
28:     seed $\leftarrow$ dequeue(Queue_AFL_unmutated)
29:     mutated_seed $\leftarrow$ AFL_generate(seed) $\triangleright$ AFL-based mutation on CPU
30:     enqueue(Queue_AFL, mutated_seed)
31:   end if
32:  end while
33:end function
A.3 Fine-tuning Prompt example
### Input: “‘Based on below hex {fuzzing_target} seed, mutate a new {fuzzing_target} seed. Make sure the example is complete and valid." {hex_seed} ### Output:
A.4 Implementations
As described in subsection 3.4, we developed an asynchronous job queue to communicate between the fuzzer and LLM. Besides, we revised AFL++ source code (e.g. afl-fuzz-bitmap.c) to incorporate LLAMAFUZZ’s functionalities, such as message queue and seeds evaluations.
Following previous work Radford et al. (2019); Wang et al. (2022), we selected a temperature of 1.0 to mutate the structured data for precise and factual responses. In addition, we adopted the model quantization Polino et al. (2018) into mixed precision floating-point $16fp$ and enabled Lora Hu et al. (2021) to freeze some of the parameters, increasing training and inference speed without sacrificing too much accuracy.
A.5 Experiment Setup and Metrics
Table 5: Targets information. The programs tested under the Magma are utilized in their default versions. For programs included in the real-world bench, we specify the exact versions used by listing the Git SHA identifiers behind each project name.
Project & version Fuzz Target File format
Magma V1.2
libpng libpng_read_fuzzer PNG
libsndfile sndfile_fuzzer Audio
libtiff tiff_read_rgba_fuzzer TIFF
tiffcp
libxml2 read_memory_fuzzer XML
xmllint
lua Targets lua Lua
openssl asn1, asn1parse, bignum, server, client, x509 Binary blobs
php json JSON
exif EXIF
unserialize Serialize object
parser PHP
poppler pdf_fuzzer, pdfimages, pdftoppm PDF
sqlite3 sqlite3_fuzz SQL query
Real-world programs
mruby 14c2179 mruby_fuzzer mruby
php afa034d php_exec php
quickjs 91459fb eval javascript
binutils 7320840 fuzz_nm, fuzz_objcopy, fuzz_readelf ELF
fuzz_strings String
bloaty 34f4a66 fuzz_target ELF, Mach-O, WebAssembly
freetype2 cd02d35 ftfuzzer TTF, OTF, WOFF
grok b9286c2 decompress_fuzzer JPEG 2000
kamailio 3f774f3 fuzz_parse_msg sip_msg
fuzz_uri URI
libavc 828cdb7 avc_dec_fuzzer AVC
mvc_dec_fuzzer MVC
svc_dec_fuzzer SVC
libhevc d0897de hevc_dec_fuzzer HEVC
openh264 decoder_fuzzer H.264/MPEG-4
1c23887 AVC
zlib 0f51fb4 uncompress_fuzzer Zlib compressed
We chose Magma for several reasons. First, Magma involves a wide range of popular programs that process diverse structured data with real-world environments, including 9 libraries and 21 objects. Second, unlike LAVA-M Dolan-Gavitt et al. (2016), which primarily employs synthetic bugs and magic byte comparisons, Magma offers a diverse range of real vulnerabilities, with a total of 138 bugs spanning integer errors, divide-by-zero faults, memory overflows, use-after-free, double-free, and null-pointer dereference scenarios. It incorporates real-world bugs from older versions of software updated to their latest releases, ensuring the benchmark’s relevance and practical applicability. Third, Magma focuses on bug counts and time-to-bug metrics as more direct surrogates for performance Hazimeh et al. (2020). In addition, we chose FuzzBench in later real-world experiments, because it employed Docker containers to standardize the testing environment for each fuzzer. This setup guaranteed fairness that all fuzzers operated under identical conditions, thus ensuring comparability of results.
To assess the statistical significance of our results, we applied the Mann-Whitney U Test Klees et al. (2018). As per the Mann-Whitney U-test, a result is statistically significant if the p-value is less than 0.05. In addition, we used the Vargha-Delaney statistic Olsthoorn et al. (2020) to quantify effect size, providing insight into the magnitude of observed differences. The result was classified as negligible if $\hat{A}_{12}$ was less than 0.5, as medium if it was greater than 0.5 but not exceeding 0.8, and as large if it was greater than 0.8.
Appendix B More Experiment Result
Table 6: Average branch coverage achieved by LLAMAFUZZ, baseline with default initial seeds (AFL++) and baseline with default initial seeds combined with trimmed seeds set from LLAMAFUZZ’s fine-tuning dataset (AFL++).
Fuzz target Fuzz object Branch coverage (avg)
AFL++ LLAMAFUZZ integrated
llama2-7b-chat-hf LLaMA-3-8B-Instruct Mistral-7B-Instruct-v0.2 Qwen2-7B-Instruct
binutils fuzz_nm $9017$ 13 95813\,958 14 42314\,423 14 92914\,929 15 04115\,041
grok grk_decompress_fuzzer $2313$ $3750$ $4028$ $3839$ $4123$
kamailio fuzz_parse_msg $2692$ $3743$ $3821$ $3911$ $4013$
B.1 Performance on bug-triggered time
Figure 7: Heatmap of expected bug trigger time achieved by LLAMAFUZZ, AFL++, Moptafl, Honggfuzz, and Fairfuzz at the end of each 24-hour trail. Within each block, more intense blue shades denote shorter expected trigger times, while the grey parts represent bugs that were not triggered in any trials by that fuzzer.
Consequently, we listed all the unique bugs triggered, including bug ID and the expected time taken to trigger it in Figure 7. The reported time accounts for missed measurements (where the fuzzer only triggered a bug in M out of N campaigns) and fits the distribution of time-to-bug samples onto an exponential distribution Hazimeh et al. (2020).
B.2 Different models comparison
Constrained by GPU resources, we selected a small set of real-world programs to demonstrate the generalizability of our method across different LLMs. We select llama2-7b-chat-hf Touvron et al. (2023), LLaMA-3-8B-InstructDubey et al. (2024), Mistral-7B-Instruct-v0.2Jiang et al. (2023), and Qwen2-7B-InstructYang et al. (2024a) in this experiment to balance GPU cost. Table 6 outlines the results, comparing the performance of our LLM-enhanced fuzzing approaches against the baseline AFL++. Across all evaluated targets, LLM-enhanced methods outperformed AFL++, achieving higher branch coverage. Notably, more recent LLMs, such as LLaMA3-8B, Mistral, and Qwen-7B, exhibited superior performance, likely due to their improved inference capabilities and better contextual understanding. The result demonstrates the generalizability of our method, confirming its adaptability to various LLMs while consistently enhancing fuzzing effectiveness.
Appendix C Related work
C.1 Fuzzing
Fuzzing Zalewski (2016); Fioraldi et al. (2020b); Xu et al. (2024); Pham et al. (2019); Swiecki is an automated random software testing technique to discover vulnerabilities and bugs in the target programs or applications. Traditional fuzzers can be categorized into black-box fuzzers Feng et al. (2021); Pferscher and Aichernig (2022); Li et al. (2018); Liu et al. (2024); Borcherding et al. (2023); Householder and Foote (2012), white-box fuzzers Yang et al. (2024b); Borkar et al. (2024); Zhang et al. (2023); Arcuri et al. (2024), and greybox fuzzers Blackwell et al. (2025); Pham et al. (2019); Böhme et al. (2017); Kim et al. (2023); Fioraldi et al. (2020a); Huang et al. (2022); Wang et al. (2019b) depending on whether fuzzers are aware of the program structure. The black-box fuzzer threat targets a black box, and it’s unaware of the program structure. Usually, a black-box fuzzer has a high execution volume since it randomly generates test input, but it only scratches the surface. YARPGen Livinskii et al. (2020) applies random mutation rigorously applies language specifications to ensure the validity of test cases to test C and C++ compilers. Similarly, Csmith Yang et al. (2011) generates programs that cover a large subset of C while avoiding undefined and unspecified behaviors.
White-box fuzzers utilize program analysis to improve the code coverage to explore certain code regions, which can be efficient in revealing vulnerabilities in complex logic. WhisperFuzz Borkar et al. (2024) introduces a static analysis method designed specifically to detect and locate timing vulnerabilities in processors. The tool focuses on evaluating the coverage of microarchitectural timing behaviors, providing a targeted and comprehensive assessment that aids in identifying potential security risks associated with timing flaws.
However, program analysis and defining specialized seed generation grammar could be extremely time-consuming. Greybox fuzzer combines the effectiveness of white-box fuzzer and the efficiency of black-box fuzzer. It leverages instrumentation to get feedback from target programs and leading fuzzers to generate more valuable seeds resulting in higher code coverage. Greybox fuzzers usually combined with mutation strategies rely on iterative modifications of existing seeds to produce novel fuzzing inputs. In addition to basic mutations, recent researchers have developed complex transformations to maintain type consistency Jain et al. (2018); Chaliasos et al. (2022), adding historical bug-triggering code snippets Holler et al. (2012); Zhao et al. (2022), and coverage feedback Aschermann et al. (2019); Fioraldi et al. (2020b) for improved testing efficiency. American Fuzzy Lop (AFL) Zalewski (2016) and its variations Fioraldi et al. (2020b); Lyu et al. (2019); Crump et al. (2023), employ genetic algorithms with a fitness function to prioritize fuzzing inputs for further mutations aimed at enhancing coverage, concentrating on byte-level changes.
C.2 Coverage-guided greybox fuzzing
To overcome the inherent randomness challenges in fuzzing, researchers suggest using bit-map to record coverage information as feedback to more effectively guide the fuzzing process Zalewski (2016). Since vulnerabilities cannot be detected in uncovered paths, focusing on expanding the coverage of execution paths is a reasonable step toward improving the performance of fuzzing techniques.

### v4

@@ 4.2. Implementation details
We implemented LLAMAFUZZ by extending AFL++ (Fioraldi et al., 2020b) to evaluate the potential of LLMs in addressing the limitations of traditional fuzzing on structured inputs. The base model used is LLaMA 2-7B-Chat (Touvron et al., 2023), chosen for its balance between efficiency and reasoning capability. Since LLAMAFUZZ is built on top of AFL++, any performance deviation can be attributed to our LLM-driven mutation module. To mitigate seed memorization, all experimental seeds were excluded from the fine-tuning dataset.
We followed the official build procedures of OSS-Fuzz (Serebryany, 2017), Fuzz-Bench (Metzman et al., 2021), and Magma (Hazimeh et al., 2020) to prepare fuzzing targets. For OSS-Fuzz benchmarks, we used the default initial seed corpus provided by the project. For Magma benchmarks, we similarly adopted the benchmark-defined default seeds to ensure experimental consistency.
Fine-tuning LLM
System Integration
As described in subsection 3.4, we developed an asynchronous job-queue system for communication between the fuzzer and the LLM. The AFL++ source code (e.g., afl-fuzz-bitmap.c) was modified to integrate LLAMAFUZZ’s key functionalities, including the message queue, seed evaluation, and mutation invocation mechanisms. Following prior work (Radford et al., 2019; Wang et al., 2022), we used a temperature of 1.0 to ensure balanced exploration and factual mutation behavior. We further adopted mixed-precision quantization (Polino et al., 2018) (FP16) and parameter freezing via LoRA (Hu et al., 2021) to accelerate both fine-tuning and inference with minimal degradation in accuracy.
@@
Figure 3. Distribution of the average number of bugs triggered over 24 hours.
Distribution of the average number of bugs triggered over 24 hours.
@@
Figure 4. Arithmetic mean number of bugs identified for each project per trial. The black line denotes the 95% confidence interval.
Arithmetic mean number of bugs identified for each project per trial. The black line denotes the 95% confidence interval.
4.3. Main results
LLAMAFUZZ find more bugs
Figure 5. Heatmap of expected bug trigger time achieved by LLAMAFUZZ, AFL++, Moptafl, Honggfuzz, and Fairfuzz at the end of each 24-hour trail. Within each block, more intense blue shades denote shorter expected trigger times, while the grey parts represent bugs that were not triggered in any trials by that fuzzer.
Heatmap of expected bug trigger time achieved by LLAMAFUZZ, AFL++, Moptafl, Honggfuzz, and Fairfuzz at the end of each 24-hour trail. Within each block, more intense blue shades denote shorter expected trigger times, while the grey parts represent bugs that were not triggered in any trials by that fuzzer.
Performance on bug-triggered time
Consequently, we listed all the unique bugs triggered, including bug ID and the expected time taken to trigger it in Figure 5. The reported time accounts for missed measurements (where the fuzzer only triggered a bug in M out of N campaigns) and fits the distribution of time-to-bug samples onto an exponential distribution (Hazimeh et al., 2020).
@@ 4.3. Main results
Table 3. Average branch coverage achieved by our approach (LLAMAFUZZ) and the baseline (AFL++). We report the average branch coverage, p-values produced by the Mann-Whitney U Test, and the Vargha-Delaney statistics ($\hat{A}_{12}$). For assessing effect size, we use the labels -, M, and L to represent negligible, medium, and large effects, respectively.
@@
binutils fuzz_nm 13 96913\,969 $9017$ 54.9154.91% ¡0.01 L(1)
fuzz_objcopy 22 11822\,118 12 49412\,494 77.03% ¡0.01 L(1)
fuzz_readelf $6552$ $4437$ 47.67% ¡0.01 L(1)
fuzz_strings $6441$ $5295$ 21.64% ¡0.01 L(1)
bloaty fuzz_target $5972$ $5722$ 4.37% ¡0.01 L(1)
freetype2 freetype2-ftfuzzer 10 52110\,521 $9978$ 5.45% ¡0.01 L(1)
grok grk_decompress_fuzzer $3721$ $2313$ 60.87% ¡0.01 L(1)
@@
4.4. Performance comparison on OSS-Fuzz
@@
First, we evaluated the performance of LLAMAFUZZ against specialized grammar-based fuzzers, including Gramatron (Srivastava and Payer, 2021), Grimoire (Blazytko et al., 2019), and Nautilus (Aschermann et al., 2019). We selected a common set of fuzzing targets compatible with all specialized grammar-based fuzzers, including mruby, PHP, and quickjs. Next, we conducted a more comprehensive evaluation using real-world programs. Given that grammar-based fuzzers are limited to well-defined, specialized targets, AFL++ was chosen as the reference competitor, representing the state-of-the-art in greybox fuzzing subsection 4.3.
@@
Figure 6. The distribution of reached bug achieved by LLAMAFUZZ, AFL++, Gramatron, Grimoire, and Nautilus.
The distribution of reached bug achieved by LLAMAFUZZ, AFL++, Gramatron, Grimoire, and Nautilus.
@@
Table 4. Average branch coverage achieved by LLAMAFUZZ, baseline with default initial seeds (AFL++) and baseline with default initial seeds combined with trimmed seeds set from LLAMAFUZZ’s fine-tuning dataset (AFL++).
Fuzz target Fuzz object Branch coverage (avg)
AFL++ LLAMAFUZZ integrated
Llama2 Llama3 Mistral Qwen2
binutils fuzz_nm $9017$ 13 95813\,958 14 42314\,423 14 92914\,929 15 04115\,041
grok grk_decompress_fuzzer $2313$ $3750$ $4028$ $3839$ $4123$
kamailio fuzz_parse_msg $2692$ $3743$ $3821$ $3911$ $4013$
Table 5. Average branch coverage achieved by LLAMAFUZZ compared with without finetuning.
Fuzz target Fuzz object Branch coverage (avg)
Llama2 Llama2 Llama3 Llama3
(w/o finetune) (w/o finetune)
binutils fuzz_nm 10 24710\,247 13 95813\,958 12 71112\,711 14 42314\,423
grok grk_decompress_fuzzer $3023$ $3750$ $3761$ $4028$
kamailio fuzz_parse_msg $3491$ $3743$ $3634$ $3821$
4.5. Ablation study: Generalization across different models
We selected a small set of real-world programs to demonstrate the generalization of our method across different LLMs. We select llama2-7b-chat-hf (Touvron et al., 2023), LLaMA-3-8B-Instruct(Dubey et al., 2024), Mistral-7B-Instruct-v0.2(Jiang et al., 2023), and Qwen2-7B-Instruct(Yang et al., 2024a) in this experiment to balance GPU cost. Table 4 outlines the results, comparing the performance of our LLM-enhanced fuzzing approaches against the baseline AFL++. Across all evaluated targets, LLM-enhanced methods outperformed AFL++, achieving higher branch coverage. Notably, more recent LLMs, such as LLaMA3-8B, Mistral, and Qwen-7B, exhibited superior performance, likely due to their improved inference capabilities and better contextual understanding. The result demonstrates the generalizability of our method, confirming its adaptability to various LLMs while consistently enhancing fuzzing effectiveness.
4.6. Ablation study: Effectiveness of fine-tuning
To examine whether fine-tuning enhances the fuzzing capability of LLAMAFUZZ, we compare average branch coverage before and after mdoel fine-tuning. As illustrate in Table 5, fine-tuning consistently improves coverage across all models and targets, confirming its effectiveness in adapting LLMs to fuzzing-specific objectives. These gains demonstrate that model adaptation enables LLAMAFUZZ to generate inputs more aligned with the syntactic and semantic structure of the fuzzing domain, leading to deeper code exploration. Notably, even the stronger Llama3 backbone benefits from task-specific fine-tuning, indicating that general-purpose instruction-tuning alone is insufficient for coverage-oriented fuzzing.
5. Analysis
@@
5.1. Did the LLM improve fuzzing by memorizing the fine-tuning data?
Table 6. Average branch coverage achieved by LLAMAFUZZ, baseline with default initial seeds (AFL++) and baseline with default initial seeds combined with trimmed seeds set from LLAMAFUZZ’s fine-tuning dataset (AFL++).
Fuzz target Fuzz object Branch coverage (avg)
LLAMAFUZZ AFL++ AFL++ Improv. p-value $\hat{A}_{12}$
binutils fuzz_nm 13 96913\,969 $9017$ 10 98310\,983 27.18% ¡0.01 L(1)
fuzz_objcopy 22 11822\,118 12 49412\,494 13 57513\,575 62.93% ¡0.01 L(1)
fuzz_readelf $6552$ $4437$ $5789$ 13.79% ¡0.01 L(1)
fuzz_strings $6441$ $5295$ $5220$ 23.39% ¡0.01 L(1)
bloaty fuzz_target $5972$ $5722$ $5918$ 0.90% 0.37 M(0.7)
grok grk_decompress_fuzzer $3721$ $2313$ $2629$ 41.54% 0.02 L(1)
kamailio fuzz_parse_msg $3743$ $2692$ $2521$ 48.49% ¡0.01 L(1)
libavc avc_dec_fuzzer $9872$ $9838$ $9847$ 0.26% 0.01 L(1)
@@ 5.1. Did the LLM improve fuzzing by memorizing the fine-tuning data?
Figure 7. Code coverage growth of LLAMAFUZZ and AFL++ on the binutils-nm target over time. Black triangles mark LLM-generated seeds, while the red vertical line highlights seeds directly sourced from LLM. The green line represents LLAMAFUZZ’s coverage, and the grey background indicates AFL++’s coverage range.
Code coverage growth of LLAMAFUZZ and AFL++ on the binutils-nm target over time. Black triangles mark LLM-generated seeds, while the red vertical line highlights seeds directly sourced from LLM. The green line represents LLAMAFUZZ’s coverage, and the grey background indicates AFL++’s coverage range.
Table 7. Byte-level comparison of an original PNG seed and its AFL- and LLAMAFUZZ-mutated variants.
@@ 5.3. Case study
6. Related Work
6.1. Fuzzing
White-box fuzzers utilize program analysis to improve the code coverage to explore certain code regions, which can be efficient in revealing vulnerabilities in complex logic. WhisperFuzz (Borkar et al., 2024) introduces a static analysis method designed specifically to detect and locate timing vulnerabilities in processors. The tool focuses on evaluating the coverage of microarchitectural timing behaviors, providing a targeted and comprehensive assessment that aids in identifying potential security risks associated with timing flaws.
However, program analysis and defining specialized seed generation grammar could be extremely time-consuming. Greybox fuzzer combines the effectiveness of white-box fuzzer and the efficiency of black-box fuzzer. It leverages instrumentation to get feedback from target programs and leading fuzzers to generate more valuable seeds resulting in higher code coverage. Greybox fuzzers usually combined with mutation strategies rely on iterative modifications of existing seeds to produce novel fuzzing inputs. In addition to basic mutations, recent researchers have developed complex transformations to maintain type consistency (Jain et al., 2018; Chaliasos et al., 2022), adding historical bug-triggering code snippets (Holler et al., 2012; Zhao et al., 2022), and coverage feedback (Aschermann et al., 2019; Fioraldi et al., 2020b) for improved testing efficiency. American Fuzzy Lop (AFL) (Zalewski, 2016) and its variations (Fioraldi et al., 2020b; Lyu et al., 2019; Crump et al., 2023), employ genetic algorithms with a fitness function to prioritize fuzzing inputs for further mutations aimed at enhancing coverage, concentrating on byte-level changes.
6.2. Coverage-guided greybox fuzzing
To overcome the inherent randomness challenges in fuzzing, researchers suggest using bit-map to record coverage information as feedback to more effectively guide the fuzzing process (Zalewski, 2016). Since vulnerabilities cannot be detected in uncovered paths, focusing on expanding the coverage of execution paths is a reasonable step toward improving the performance of fuzzing techniques.